OpenStudy (anonymous):

is this serie convergent? sigma [(n-1)/(n+1)]^n by which tests ?

OpenStudy (allank):

Hmm...I think the series diverges...lemme write it out..

OpenStudy (allank):

Trying the n'th term test, let's directly observe the stuff within the ()^n So we have limit as n approaches infinity of (n-1)/(n+1) which is 1. By intuition, we can see that limit as n->infinity [(n-1)/(n+1)]^n is not = 0. Hence the series diverges.

OpenStudy (anonymous):

Given a series: \[\sum_k a_k\] You know it DIVERGES (DOES NOT CONVERGE) if: \[\lim_{k \rightarrow \infty} a_k = 0\] This does NOT mean it converges, but it is a necessary condition for convergence. Therefore we have: \[\lim_{k \rightarrow \infty} \left( \frac{k-1}{k+1} \right)^k\] is indeterminate: \[\left(\frac{\infty}{\infty} \right)^{\infty}\] But 'eyeballing' it will show that the inside goes to 1 and an infinite product of 1's is 1 therefore it must diverge.

OpenStudy (anonymous):

thanks ..